how many seconds after the cannonball was shot will it first be 1000 feet high? and how many seconds will the cannonball next be 1000 feet? and how did you solve this?
scenario:suppose a cannon shoots a cannonball into the air so that the initial upward velocity of the cannonball is 320 feet per second.
formula: H=-1/2*gt^+vt+h where g is the acceleration due to gravity, and on the earth, equals 32ft/sec^, v is the initial upward velocity of the object and h is the initial height above the ground.

Answer by stanbon(75887)

how many seconds after the cannonball was shot will it first be 1000 feet high? and how many seconds will the cannonball next be 1000 feet? and how did you solve this?
scenario:suppose a cannon shoots a cannonball into the air so that the initial upward velocity of the cannonball is 320 feet per second.
formula: H=-1/2*gt^2+vt+h where g is the acceleration due to gravity, and on the earth, equals 32ft/sec^, v is the initial upward velocity of the object and h is the initial height above the ground.
------------
H(t)=(-1/2)(32)t^2 + 320t + 0
Let H(t) = 1000 and solve for "t".
-16t^2+320t-1000=0
-4t^2+80-250=0
t=[-80+-sqrt(80^2-4*-4*-250)]/-8
t=3.876... seconds or t=16.1237... seconds
Cheers,
Stan H.